OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x4 + 4x2

OpenStudy (mathstudent55):

Can you factor a common factor out?

OpenStudy (anonymous):

x

OpenStudy (mathstudent55):

Yes, but you can factor more than that.

OpenStudy (anonymous):

x^2?

OpenStudy (mathstudent55):

Correct. Start by factoring out x^2; what do you get?

OpenStudy (anonymous):

x^2(x^2-4) right?

OpenStudy (mathstudent55):

Wasn't the original polynomial f(x) = x4 + 4x2? (with a + between the terms?)

OpenStudy (anonymous):

oh yes i meant +

OpenStudy (mathstudent55):

We're here, then. f(x) = x^2(x^2 + 4)

OpenStudy (anonymous):

yes what do i do from there

OpenStudy (mathstudent55):

We can write the first two terms as linear terms: \(f(x) = x \cdot x \cdot (x^2 + 4) \)

OpenStudy (mathstudent55):

Now we need to deal with \(x^2 + 4\) and make it the product of two linear terms.

OpenStudy (mathstudent55):

If we had \(x^2 - 9\), can you factor this?

OpenStudy (anonymous):

(x+3)(x-3)

OpenStudy (mathstudent55):

Correct. Now let's write \(x^2 + 4\) as a difference of squares. Then you'll be able to do the same kind of factoring. Notice that \(x^2 - 9 = (x + \sqrt 9) (x - \sqrt 9) = (x + 3)(x - 3) \) Now let's do the same to this. Use the square root as I did above in the middle step. \(x^2 - (-4)\)

OpenStudy (anonymous):

so (x+squareroot of 2)(x-squareroot of 2)

OpenStudy (mathstudent55):

You're on the right track, but you need to be careful. Let's follow the example carefully. \(x^2 - (-4) = (x + \sqrt{-4} )(x - \sqrt{-4} )\) Now we deal with the imaginary roots. \(= (x + i\sqrt{4} )(x - i\sqrt{4} )\) Now we deal with \(\sqrt 4\) \(= (x + 2 i )(x - 2i)\)

OpenStudy (anonymous):

so would this be the answer f(x) = x^2 (x + 2i)(x - 2i)

OpenStudy (mathstudent55):

Yes. Unless you want to split up the x's and show it as \(f(x) = x \cdot x \cdot (x + 2i)(x - 2i) \) Now you are really showing every linear term.

OpenStudy (anonymous):

Ok thank you! i understand now.

OpenStudy (mathstudent55):

Great. You're welcome.